OpenStudy (lilai3):

Number Theory: The sum of a 2-digit number and the 2-digit number when the digits are reversed is 77. If the difference of the same two numbers is 45, what are the two 2-digit numbers?

OpenStudy (anonymous):

Let x represent the first digit and y the second. Because of the way the decimal system works, the number can be expressed as: 10^1 * x + 10^0 * y = 10x + y Then, the number reversed is: 10^1 * y + 10^0 * x = 10y + x Add those two together and set it to 77, then subtract the two and set it to 45. Solve.

OpenStudy (lilai3):

Um, I got 61 and 16?

OpenStudy (anonymous):

61 + 16 = 77 61 - 16 = 45 Seems good.

OpenStudy (lilai3):

thank you! <333 :3